Find the present value of the following amount.
$50,000 in 5 years and 6 months at 12% compounded quarterly

Answer by rfer(16322)

p=a/(1+r/n)^nt
p=50000/1+.12/4)^4*5
p=50000/(1.03)^20
p=50000/1.81
p=$27,624.31